what is the height of a crate that is 72 cubic feet and is 6 feet wide 3 feet long?


Answer by mananth(16949)

what is the height of a crate that is 72 cubic feet and is 6 feet wide 3 feet long?
The volume of a cuboid is given by L*W*H where L = length, W= width and H = height
Let the height of the crate be x
72 cu. feet = 6*3*x
72 cu, feet = 18x
x= 72/18
x= 4. The height of the crate = 4 feet.